Find the measure of an angle whose supplement is 4 times the measure of its compliment
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52787)

.

x = angle

complementary angle  90-x degrees

supplementary angle 180-x degrees


Equation

    180-x = 4*(90-x).

Solution

    180-x = 360 - 4x

    4x - x = 360 - 180

    3x     = 180

     x     = 180/3 = 60 degrees.    ANSWER

Solved.



Answer by Alan3354(69443)

Find the measure of an angle whose supplement is 4 times the measure of its compliment
---------------------
complement.
not compliment.